OpenStudy (anonymous):

This is an 11 part Q. 1. In this problem, we analyze the profit found for sales of decorative tiles. A demand equation (sometimes called a demand curve) shows how much money people would pay for a product depending on how much of that product is available on the open market. Often, the demand equation is found empirically (through experiment, or market research).

OpenStudy (anonymous):

a. Suppose a market research company finds that at a price of p = $20, they would sell x = 42 tiles each month. If they lower the price to p = $10, then more people would purchase the tile, and they can expect to sell x = 52 tiles in a month

OpenStudy (anonymous):

(q,p) = (42, 20) and (52, 10) so you will do (52 - 42) / (20 - 10) = 1 p + 52 = 1(q - 10) p = 1q - 10

OpenStudy (anonymous):

A company’s revenue is the amount of money that comes in from sales, before business costs are subtracted. For a single product, you can find the revenue by multiplying the quantity of the product sold, x, by the demand equation, p. b. Substitute the result you found from part a. into the equation R = xp to find the revenue equation. Provide your answer in simplified form.

OpenStudy (lilg132):

im not too sure but if i understand correctly Revenue = x tiles * Price so 20*42 = 840 10*52 = 520

OpenStudy (anonymous):

The costs of doing business for a company can be found by adding fixed costs, such as rent, insurance, and wages, and variable costs, which are the costs to purchase the product you are selling. The portion of the company’s fixed costs allotted to this product is $300, and the supplier’s cost for a set of tile is $6 each. Let x represent the number of tile sets. c. If b represents a fixed cost, what value would represent b? d. Find the cost equation for the tile. Write your answer in the form C = mx + b.